OpenStudy (anonymous):

True Or False The product of two sine functions equals the sine of the product. What the heck does this mean???

OpenStudy (anonymous):

I just cant read it. 13

Parth (parthkohli):

Is this an identity?\[\sin(x)\sin(y) = \sin(xy)\]

jimthompson5910 (jim_thompson5910):

I think it's asking if sin(A) * sin(B) = sin(A*B) is true

jimthompson5910 (jim_thompson5910):

yeah what ParthKohli posted as well (not 100% sure though)

OpenStudy (anonymous):

I dont think so. My calculator says it is not.

jimthompson5910 (jim_thompson5910):

if you can find/show a counter-example, then it's false

OpenStudy (anonymous):

ok, so for example sin(5)^2 /= sin(25) so its false

jimthompson5910 (jim_thompson5910):

that counter-example works